# Solver-Free Decision-Focused Learning for Linear Optimization Problems

### v1

@@ 5.1 Experimental setup
Here we give an overview of our experimental setup; more details can be found in the appendix. The models are linear regressors, as is common in existing literature spo ; ltr ; dfl_survey ; NCE ; schutte2023robust . We train these models using the Adam optimizer adam . All hyperparameters are tuned on independent validation sets prior to training. We train the models until their performance on the validation set has not improved by at least $1\%$ for three checks in a row, or until training time has surpassed $600$ seconds (not including evaluation on the validation set), whichever comes first. For lava, this training time includes both adjacent vertex precomputation and actual training. We report the test set performance of the model state that led to the best validation set performance during training. When reporting training times, we report the time it took to reach this state. Reported results are the average taken over $5$ independent runs with varying train-validation-test splits. We report the normalized regret:
Normalized $\textit{Normalized Regret}=\frac{\sum_{i=1}^{n_{test}}c^{\top}z^{\star}(\hat{c})-c^{\top}z^{\star}(c)}{\sum_{i=1}^{n_{test}}c^{\top}z^{\star}(\hat{c})}$ (5)
All experiments were conducted on a machine equipped with an Intel(R) Core(TM) i7-1165G7 processor and 16 GB of RAM. All code is implemented in Python, and builds on the PyEPO library pyepo . All code and data will be made available upon acceptance of the paper.
@@ 5.3 Benchmarks
We compare these methods on the following predict-then-optimize benchmarks, listed in increasing order of degeneracy. We refer to the appendix for further details.
@@
    Random LPs: We construct LPs by sampling the entries of $A$ and $b$ uniformly at random, which ensures that the feasible space contains only nondegenerate vertices gonzaga2007generation . Each LP contains 150 variables and 50 constraints. We make sure that all generated constraints are nonredundant. As true mapping between features and cost values, we use a degree 8 polynomial that is approximated using linear regression to simulate model misspecification. This is common practice in DFL evaluations spo ; ltr ; pyepo ; tang2024cave ; schutte2023robust , and is especially relevant when training interpretable models interpretable1 ; interpretable3 ; interpretable2 .
@@ 5.4 Results and discussion
Figure 2 shows the training time and test set regret (log scale) of the various methods (full results with standard errors are given in tabular format in the appendix). lava is Pareto-efficient on all
@@
benchmarks, meaning it is never bested in both training time and test regret by another method. On the random LPs and multi-dimensional knapsack problems, it offers a highly desirable trade-off: training time is reduced significantly with minimal degradation in solution quality compared to the best-performing methods (SPO+ and PFYL). On the highly degenerate shortest path problem, the relative test regret of lava is slightly worse, and the improvement in training time is less significant. Despite the shortest path being the smallest problem considered, it is the benchmark on which lava takes the longest. Figure 3 shows why: the majority of lava’s computation time is spent precomputing the adjacent vertices, which is computationally expensive due to the shortest path’s high degree of degeneracy (see Section 4.2.2). Because of this, lava may not be the most appropriate method to address network flow problems, which tend to be highly degenerate bertsimas1997introduction . Fortunately, these problems do not suffer much from efficiency issues in the first place, as specialized solving algorithms are available linear_programming_and_network_flows ; network_simplex . Figure 3 also highlights another advantage of lava. In hyperparameter tuning, the adjacent vertices can be precomputed once and subsequently used in each configuration explored. Since the majority of lava’s computation time is spent on this precomputation, the tuning process becomes particularly efficient.
@@ 6 Conclusions and future work
The forward implication holds trivially: if $z$ is an optimal solution for cost vector $c$, then $\forall z^{\prime}\in\mathcal{F}:c^{\top}z\leq c^{\top}z^{\prime}$, where $\mathcal{F}$ denotes the feasible region (=$\mathcal{F}=\left\{z\in\mathbb{R}^{n}\;\middle|\;Az=b,\;z\geq 0\right\}$). Since $Z_{adj}(z)\subseteq\mathcal{F}$, it holds that $\forall z_{adj}\in Z_{adj}(z):c^{\top}z\leq c^{\top}z_{adj}$.
@@ Appendix B Ablation of thresholding in lava
lava ($\epsilon=0.1$) $0.016\pm 0.001$ $0.086\pm 0.002$ $0.060\pm 0.012$
lava ($\epsilon=0$) $0.023\pm 0.001$ $0.127\pm 0.007$ $0.061\pm 0.008$
lava (no threshold) $0.166\pm 0.009$ $0.156\pm 0.003$ $0.243\pm 0.022$
@@ Appendix C More information on the TNP rule
To apply TNP pivoting, we must start from a transition node associated with $z^{\star}$. If our initial basis is not a transition node, we can obtain one by applying random pivots (with respect to an anti-cycling rule [8]). Since the starting node is now a transition node, there will be a column $d$ of $D=-B^{-1}N$ such that $\min\limits_{\{i\mid d_{i}<0\}}\left(-\frac{z^{\star}_{B_{curr}(i)}}{d_{i}}\right)>0$, i.e., $\theta^{\star}>0$ in the minimum ratio test (line 11 in Algorithm 1). This column is referred to as the transition column, and is denoted by $t$. This column will remain a transition column throughout all the subsequent TNP pivots. The TNP rule comes into effect when, for an entering variable with index $j$, $\theta^{\star}=0$ (line 11 of Algorithm 1) and there are multiple minimizers, i.e., $\left|I_{min}^{(j)}\right|>1$, where $I_{min}^{(j)}$ = $\{i\mid-\frac{z^{\star}_{B_{curr}(i)}}{d_{i}}=0\}$. The TNP rule is used to decide which $i\in I_{min}$ to select as leaving variable. This $i$ is selected as $\max_{k}\{\frac{D_{kt}}{D_{kj}}|k\in I_{min}^{(j)}\}$, which ensures that $t$ remains a transition column. For a proof of this statement, as well as further details on the TNP rule, we refer the reader to [15] and [17].
@@ D.2 Optimization problems
Given a set of items =$\mathcal{I}=\{1,\ldots,n\}$, where each item $i\in\mathcal{I}$ has a value $c_{i}$ and a weight $w_{ij}$ for each of $m$ knapsack constraints, the multi-dimensional knapsack problem seeks to find the subset of items with maximum total value without exceeding the capacity $W_{j}$ for each constraint $j\in\{1,\ldots,m\}$.
@@ D.3 True predictive mappings
    where $\epsilon_{j}$ is sampled uniformly from $[1-\bar{\epsilon},1+\bar{\epsilon}]$.
@@ Appendix E Results in tabular form
CaVE $0.017\pm 0.001$ $0.111\pm 0.002$ $0.080\pm 0.014$
lava (Ours) $0.016\pm 0.001$ $0.086\pm 0.002$ $0.060\pm 0.012$

### v2

@@ 5.1 Experimental setup
Here we give an overview of our experimental setup; more details can be found in Appendix D. The models are linear regressors, as is common in existing literature spo ; ltr ; dfl_survey ; NCE ; schutte2023robust . We train these models using the Adam optimizer adam . All hyperparameters are tuned on independent validation sets prior to training. We train the models until their performance on the validation set has not improved by at least $1\%$ for three checks in a row, or until training time has surpassed $600$ seconds (not including evaluation on the validation set), whichever comes first. For lava, this training time includes both adjacent vertex precomputation and actual training. We report the test set performance of the model state that led to the best validation set performance during training. When reporting training times, we report the time it took to reach this state. Reported results are the average taken over $5$ independent runs with varying train-validation-test splits. We report the normalized regret:
$\textit{Normalized Regret}=\frac{\sum_{i=1}^{n_{test}}c^{\top}z^{\star}(\hat{c})-c^{\top}z^{\star}(c)}{\sum_{i=1}^{n_{test}}c^{\top}z^{\star}(\hat{c})}$ (5)
All experiments were conducted on a machine equipped with an Intel(R) Core(TM) i7-1165G7 processor and 16 GB of RAM. All code is implemented in Python, and builds on the PyEPO library pyepo . Our code and data are available at https://github.com/ML-KULeuven/Solver-Free-DFL/.
@@ 5.3 Benchmarks
We compare these methods on the following predict-then-optimize benchmarks, listed in increasing order of degeneracy. We refer to Appendix D for further details.
@@
    Random LPs: We construct LPs by sampling the entries of $A$ and $b$ uniformly at random, which ensures that the feasible space contains only nondegenerate vertices gonzaga2007generation . Each LP contains 150 variables and 50 constraints. We make sure that all generated constraints are nonredundant. As true mapping between features and cost values, we use a degree 8 polynomial that is approximated using linear regression to simulate model misspecification. This is common practice in DFL evaluations spo ; ltr ; schutte2023robust ; pyepo ; tang2024cave , and is especially relevant when training interpretable models interpretable1 ; interpretable3 ; interpretable2 .
@@ 5.4 Results and discussion
Figure 2 shows the training time and test set regret (log scale) of the various methods (full results with standard errors are given in tabular format in Appendix E). lava is Pareto-efficient on all benchmarks, meaning it is never bested in both training time and test regret by another method. On the random LPs and multi-dimensional knapsack problems, it offers a highly desirable trade-off: training time is reduced significantly with minimal degradation in solution quality compared to the best-performing methods (SPO+ and PFYL). On the highly degenerate shortest path problem, the relative test regret of lava is slightly worse, and the improvement in training time is less significant. Despite the shortest path being the smallest problem considered, it is the benchmark on which lava takes the longest.
@@
Figure 3 shows why: the majority of lava’s computation time is spent precomputing the adjacent vertices, which is computationally expensive due to the shortest path’s high degree of degeneracy (see Section 4.2.2). Because of this, lava may not be the most appropriate method to address network flow problems, which tend to be highly degenerate bertsimas1997introduction . Fortunately, these problems do not suffer much from efficiency issues in the first place, as specialized solving algorithms are available linear_programming_and_network_flows ; network_simplex . Figure 3 also highlights another advantage of lava. In hyperparameter tuning, the adjacent vertices can be precomputed once and subsequently used in each configuration explored. Since the majority of lava’s computation time is spent on this precomputation, the tuning process becomes particularly efficient.
@@ 6 Conclusions and future work
7 Acknowledgements
This research received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (Grant No. 101002802, CHAT-Opt). Senne Berden is a fellow of the Research Foundation-Flanders (FWO-Vlaanderen, 11PQ024N).
@@ 7 Acknowledgements
The forward implication holds trivially: if $z$ is an optimal solution for cost vector $c$, then $\forall z^{\prime}\in\mathcal{F}:c^{\top}z\leq c^{\top}z^{\prime}$, where $\mathcal{F}$ denotes the feasible region ($\mathcal{F}=\left\{z\in\mathbb{R}^{n}\;\middle|\;Az=b,\;z\geq 0\right\}$). Since $Z_{adj}(z)\subseteq\mathcal{F}$, it holds that $\forall z_{adj}\in Z_{adj}(z):c^{\top}z\leq c^{\top}z_{adj}$.
@@ Appendix B Ablation of thresholding in lava
In Table 1, we show the importance of thresholding the lava loss at zero (i.e., $\epsilon=0$ or using a small constant (e.g., $\epsilon=0.1$), in comparison with removing the threshold (i.e., $\epsilon=\infty$). In Figure 5, we show the effect of using a small non-zero value for $\epsilon$ on the learning curves, using the random LP benchmark.
@@
lava ($\epsilon=0$) $0.023\pm 0.001$ $127\pm 0.007 0.0$$061\pm 0.008$
lava ($\epsilon=0.1$) $016\pm 0.001 0.0$$86\pm 0.002 0.0$$060\pm 0.012$
lava ($\epsilon=\infty$) $0.166\pm 0.009$ $0.156\pm 0.003$ $0.243\pm 0.022$
Figure 5: Learning curves for lava on the random LP benchmark, comparing margin parameter values $\epsilon=0$ and $\epsilon=0.1$.
@@ Appendix C More information on the TNP rule
To apply TNP pivoting, we must start from a transition node associated with $z^{\star}$. If our initial basis is not a transition node, we can obtain one by applying random pivots (with respect to an anti-cycling rule [8]). Since the starting node is now a transition node, there will be a column $d$ of $D=-B^{-1}N$ such that $\min\limits_{\{i\mid d_{i}<0\}}\left(-\frac{z^{\star}_{B_{curr}(i)}}{d_{i}}\right)>0$, i.e., $\theta^{\star}>0$ in the minimum ratio test (line 13 in Algorithm 1). This column is referred to as the transition column, and is denoted by $t$. This column will remain a transition column throughout all the subsequent TNP pivots. The TNP rule comes into effect when, for an entering variable with index $j$, $\theta^{\star}=0$ (line 13 of Algorithm 1) and there are multiple minimizers, i.e., $\left|I_{min}^{(j)}\right|>1$, where $I_{min}^{(j)}$ = $\{i\mid-\frac{z^{\star}_{B_{curr}(i)}}{d_{i}}=0\}$. The TNP rule is used to decide which $i\in I_{min}$ to select as leaving variable. This $i$ is selected as $\max_{k}\{\frac{D_{kt}}{D_{kj}}|k\in I_{min}^{(j)}\}$, which ensures that $t$ remains a transition column. For a proof of this statement, as well as further details on the TNP rule, we refer the reader to [15] and [17].
@@ D.2 Optimization problems
Given a set of items $\mathcal{I}=\{1,\ldots,n\}$, where each item $i\in\mathcal{I}$ has a value $c_{i}$ and a weight $w_{ij}$ for each of $m$ knapsack constraints, the multi-dimensional knapsack problem seeks to find the subset of items with maximum total value without exceeding the capacity $W_{j}$ for each constraint $j\in\{1,\ldots,m\}$.
@@ D.3 True predictive mappings
    where $\epsilon_{j}$ is sampled uniformly from $1-\bar{\epsilon},1+\bar{\epsilon}]$.
@@ Appendix E Results in tabular form
Tables 2 and 3 show the results of Figure 2 in tabular form.
@@
CaVE $0.017\pm 0.001$ $111\pm 0.002 0.0$$080\pm 0.014$
lava (Ours) $016\pm 0.001 0.0$$86\pm 0.002 0.0$$060\pm 0.012$